# Rank-One Modified Value Iteration

### v2

@@ Abstract.
1Delft Center for Systems and Control, Delft University of Technology, The Netherlands. 2Department of Mechanical and Industrial Engineering, University of Toronto, Canada. Correspondence to: Arman S. Kolarijani $<$a.sharifikolarijani@tudelft.nl$>$.
@@ 1. Introduction
Notations. The set of real numbers is denoted by $\mathbb{R}$. For a vector $\bm{v}\in\mathbb{R}^{n}$, we use $\bm{v}(i)$ and $[\bm{v}](i)$ to denote its $i$-th element. Similarly, $\bm{M}(i,j)$ and $[\bm{M}](i,j)$ denote the element in $i$-th row and $j$-th column of the matrix $\bm{M}\in\mathbb{R}^{m\times n}$. We use $\left\langle\bm{v},\bm{u}\right\rangle=\sum_{i=1}^{n}\bm{v}(i)\cdot\bm{u}(i)$ to denote the the inner product of the two vectors $\bm{v},\bm{u}\in\mathbb{R}^{n}$. $\left\|\bm{v}\right\|_{1}=\sum_{i=1}^{n}|\bm{v}(i)|$, $\left\|\bm{v}\right\|_{2}=\sqrt{\left\langle\bm{v},\bm{v}\right\rangle}$, and $\left\|\bm{v}\right\|_{\infty}=\max_{i=1}^{n}|\bm{v}(i)|$ denote the 1-norm, 2-norm, and $\infty$-norm of the vector $\bm{v}\in\mathbb{R}^{n}$, respectively. We use $\rho(\bm{M})$ to denote the spectral radius (i.e., the largest eigenvalue in absolute value) of a square matrix $\bm{M}\in\mathbb{R}^{n\times n}$. Given a set $\mathcal{X}$, $\Delta(\mathcal{X})$ denotes the set of probability distributions on $\mathcal{X}$. Let $x\sim P$ be a random variable with distribution $P\in\Delta(\mathcal{X})$. We use $\widehat{x}\sim P$ to denote a sample of the random variable $x$ drawn from the sample space $\mathcal{X}$ of $x$ according to the distribution $P$. We use $\bm{1}$, $\bm{0}$, and $\bm{I}$ to denote the all-one vector, the all-zero vector, and the identity matrix, respectively, with their dimension being clear from the context.
@@ 2. Optimal control of MDPs
$[\mathbf{T}(\bm{v}^{\star})](s)\coloneqq\min_{a\in\mathcal{A}}\left\{\bm{c}(s,a)+\gamma\ \mathds{E}_{s^{+}\sim P(\cdot|s,a)}\left[\bm{v}^{\star}(s^{+})\right]\right\},\quad\forall s\in\mathcal{S}.$ (3)
@@
intialize $\begin{array}[]{l}\texttt{intialize }\bm{v}_{0}\in\mathbb{R}^{|\mathcal{S}|}=\mathbb{R}^{n}\\ \texttt{for }k=0,1,\ldots\\ \hskip 14.22636pt\bm{v}_{k+1}(s)=[\mathbf{T}(\bm{v}_{k})](s),\quad\forall s\in\mathcal{S}\\ \texttt{endfor }\end{array}$ (4)
@@
$\begin{array}[]{l}\texttt{intialize }\bm{q}_{0}\in\mathbb{R}^{|\mathcal{S}\times\mathcal{A}|}=\mathbb{R}^{nm}\\ \texttt{for }k=0,1,\ldots\\ \hskip 11.38092pt\texttt{for }(s,a)\in\mathcal{S}\times\mathcal{A}\\ \hskip 22.76228pt\widehat{s}_{k}^{~{}+}\sim P(\cdot|s,a)\\ \hskip 22.76228pt\delta_{k}=\bm{q}_{k}(s,a)-[\widehat{\mathbf{T}}(\bm{q}_{k},\widehat{s}_{k}^{~{}+})](s,a)\\ \hskip 22.76228pt\bm{q}_{k+1}(s,a)=\bm{q}_{k}(s,a)-\lambda_{k}\delta_{k}\\ \hskip 11.38092pt\texttt{endfor }\\ \texttt{endfor }\end{array}$ (5)
@@ 4. Rank-one Q-learning (R1-QL)
where $\lambda_{k}\geq 0$ are properly chosen step-sizes satisfying the Robbins–Monro conditions (e.g., $\lambda_{k}=1/(k+1)$), and $\widehat{\mathbf{T}}_{k}$ is the empirical Bellman operator evaluated at iteration $k$, i.e., $[\widehat{\mathbf{T}}_{k}(\bm{q}_{k})](s,a)\coloneqq[\widehat{\mathbf{T}}(\bm{q}_{k},\widehat{s}_{k}^{~{}+})](s,a)$ with $\widehat{s}_{k}^{~{}+}\sim P(\cdot|s,a)$ for each $(s,a)\in\mathcal{S}\times\mathcal{A}$ – the subscript $k$ in $\widehat{\mathbf{T}}_{k}$ denotes the dependence on the next-state sample $\widehat{s}_{k}^{~{}+}$ generated at iteration $k$.
@@
Now, using the approximation $\widehat{\bm{d}}_{k-1}\approx\widehat{\bm{P}}_{k-1}^{\top}\widehat{\bm{d}}_{k-1}$ (i.e., assuming $\widehat{\bm{d}}_{k-1}$ is a stationary distribution of $\widehat{\bm{P}}_{k-1}$ which does not hold exactly since $\widehat{\bm{d}}_{k-1}$ is only an approximation of a stationary distribution of $\widehat{\bm{P}}_{k-1}$), we derive,
$\widehat{\bm{d}}_{k}=(1-\lambda_{k})\widehat{\bm{d}}_{k-1}+\lambda_{k}\bm{F}_{k}^{\top}\widehat{\bm{d}}_{k-1},$
or, equivalently,
$\widehat{\bm{d}}_{k}(s,a)=(1-\lambda_{k})\widehat{\bm{d}}_{k-1}(s,a)+\lambda_{k}\widehat{\bm{d}}_{k-1}(s^{+},a^{+}),\quad\forall(s,a)\in\mathcal{S}\times\mathcal{A},$ (18)
where $s^{+}$ and $a^{+}$ are given in (17). Observe that the approximation $\widehat{\bm{d}}_{k-1}\approx\widehat{\bm{P}}_{k-1}^{\top}\widehat{\bm{d}}_{k-1}$ significantly reduces the memory and time complexity of the algorithm since we do not need to keep track of the estimates $\widehat{\bm{P}}_{k}$ of the state-action transition probability matrix and perform full matrix-vector multiplications for updating the estimates $\widehat{\bm{d}}_{k}$ of the stationary distribution.
@@
8:        $\bm{f}(s,a)=(1-\lambda_{k})\widehat{\bm{d}}_{k-1}(s,a)+\lambda_{k}\widehat{\bm{d}}_{k-1}(\widehat{s}^{~{}+}_{k},\widehat{a}_{k}^{~{}+})$;
@@
10:     $\widehat{\bm{d}}_{k}=\bm{f}/\left\|\bm{f}\right\|_{1}$;
@@
The complete description of the proposed R1-QL algorithm is provided in Algorithm 2. We again note that the vector $\bm{f}$ and its normalization are only introduced to avoid the accumulation of numerical errors. The following result discusses the convergence of the proposed algorithm.
@@ 6. Limitations and future research directions
where $e_{s}$ is the ${s}$-th unit vector and $\mathbf{T}$ is now the undiscounted Bellman operator. Now, observe that
@@ Appendix A Technical Proofs
[…]
@@
$&\bm{v}_{k+1}=\mathbf{T}^{(k+1)}(\bm{v}_{0})+\beta_{k+1}\bm{1}\\ &\beta_{k+1}=\gamma\beta_{k}+\alpha_{k},\quad\text{\emph{with}}~{}\beta_{0}=0.\end{aligned}\right.$
@@
[…]
@@
$&\bm{q}_{k+1}=\bm{q}^{\emph{\text{QL}}}_{k+1}+\beta_{k+1}\bm{1}\\ &\beta_{k+1}=(1-\lambda_{k})\beta_{k}+\gamma\lambda_{k}\beta_{k}+\alpha_{k},\quad\text{\emph{with}}~{}\beta_{0}=0.\end{aligned}\right.$ (26)
@@
=(24)${=}(1-\lambda_{k})(\bm{q}^{\text{QL}}_{k}+\beta_{k}\bm{1})+\lambda_{k}\big{(}\widehat{\mathbf{T}}_{k}(\bm{q}^{\text{QL}}_{k})+\gamma\beta_{k}\bm{1}\big{)}+\alpha_{k}\bm{1}$
@@
is a scaled, weighted average of Cesaro means of the sequences $[\overline{\mathbf{T}}_{k}(\bm{q}^{\text{QL}}_{k})-\bm{q}^{\text{QL}}_{k}](s,a)$ that converge to zero almost surely for each $(s,a)\in\mathcal{S}\times\mathcal{A}$ (recall that $\lambda_{k}=1/(k+1)$ and $\bm{q}^{\text{QL}}_{k}\rightarrow\bm{q}^{\star}=\overline{\mathbf{T}}(\bm{q}^{\star})$ almost surely). ∎
@@ B.1. Algorithms
[…]
@@
    $&\bm{z}_{k}=\bm{v}_{k}-\bm{v}_{k-1},\\ &\bm{z^{\prime}}_{k}=\mathbf{T}(\bm{v}_{k})-\mathbf{T}(\bm{v}_{k-1}),\\ &\delta_{k}=\begin{cases}0,&\bm{z}_{k}^{\top}(\bm{z}_{k}-\bm{z}_{k}^{\prime})=0,\\ \frac{\bm{z}_{k}^{\top}\big{(}\bm{v}_{k}-\mathbf{T}(\bm{v}_{k})\big{)}}{\bm{z}_{k}^{\top}(\bm{z}_{k}-\bm{z}_{k}^{\prime})},&\text{otherwise},\end{cases}\\ &\bm{v}_{k+1}=(1-\delta_{k})\mathbf{T}(\bm{v}_{k})+\delta_{k}\mathbf{T}(\bm{v}_{k-1}).\end{aligned}$
@@
    $&\texttt{for }(s,a)\in\mathcal{S}\times\mathcal{A}\\ &\qquad\widehat{s}^{~{}+}\sim P(\cdot|s,a),\\ &\qquad\bm{z}_{k}(s,a)=\bm{c}(s,a)+\gamma\min_{a^{+}\in\mathcal{A}}\bm{q}_{k}(\widehat{s}^{~{}+},a^{+}),\\ &\qquad\bm{z}_{k}^{\prime}(s,a)=\bm{c}(s,a)+\gamma\min_{a^{+}\in\mathcal{A}}\bm{q}_{k-1}(\widehat{s}^{~{}+},a^{+}),\\ &\texttt{endfor }\\ &\qquad\bm{q}_{k+1}=\bm{q}_{k}+\frac{1}{1+k}(\bm{z}_{k}^{\prime}-\bm{q}_{k})+\frac{k}{1+k}(\bm{z}_{k}-\bm{z}_{k}^{\prime}).\end{aligned}$
@@
    $&\bm{F}_{k}=\bm{0}\in\mathbb{R}^{|\mathcal{S}\times\mathcal{A}|\times|\mathcal{S}\times\mathcal{A}|},\\ &\texttt{for }(s,a)\in\mathcal{S}\times\mathcal{A}\\ &\qquad\widehat{s}^{~{}+}\sim P(\cdot|s,a),\\ &\qquad\widehat{a}^{~{}+}=\operatorname*{\arg\!\min}_{a^{+}\in\mathcal{A}}\bm{q}_{k}(\widehat{s}^{~{}+},a^{+}),\\ &\qquad\big{[}\widehat{\mathbf{T}}_{k}(\bm{q}_{k})\big{]}(s,a)=\bm{c}(s,a)+\gamma\min_{a^{+}\in\mathcal{A}}\bm{q}_{k}(\widehat{s}^{~{}+},a^{+})-\bm{q}_{k}(s,a),\\ &\qquad\bm{F}_{k}\big{(}(s,a),(\widehat{s}^{~{}+},\widehat{a}^{~{}+})\big{)}=1,\\ &\texttt{endfor }\\ &\widehat{\bm{P}}_{k}=\widehat{\bm{P}}_{k-1}+\frac{1}{2+k}(\bm{F}_{k}-\widehat{\bm{P}}_{k-1}),\\ &\bm{q}_{k+1}=\bm{q}_{k}+\frac{1}{1+k}(\bm{I}-\gamma\widehat{\bm{P}}_{k})^{-1}\big{(}\widehat{\mathbf{T}}_{k}(\bm{q}_{k})-\bm{q}_{k}\big{)}.\end{aligned}$

### v3

@@ Abstract.
1Delft Center for Systems and Control, Delft University of Technology, The Netherlands. 2Department of Mechanical and Industrial Engineering, University of Toronto, Canada. Correspondence to: Arman S. Kolarijani <<a.sharifikolarijani@tudelft.nl>>.
@@ 1. Introduction
Notations. The set of real numbers is denoted by $\mathbb{R}$. For a vector $\bm{v}\in\mathbb{R}^{n}$, we use $\bm{v}(i)$ and $\bm{v}](i)$ to denote its $i$-th element. Similarly, $\bm{M}(i,j)$ and $\bm{M}](i,j)$ denote the element in $i$-th row and $j$-th column of the matrix $\bm{M}\in\mathbb{R}^{m\times n}$. We use $\left\langle\bm{v},\bm{u}\right\rangle=\sum_{i=1}^{n}\bm{v}(i)\cdot\bm{u}(i)$ to denote the the inner product of the two vectors $\bm{v},\bm{u}\in\mathbb{R}^{n}$. $\left\|\bm{v}\right\|_{1}=\sum_{i=1}^{n}|\bm{v}(i)|$, $\left\|\bm{v}\right\|_{2}=\sqrt{\left\langle\bm{v},\bm{v}\right\rangle}$, and $\left\|\bm{v}\right\|_{\infty}=\max_{i=1}^{n}|\bm{v}(i)|$ denote the 1-norm, 2-norm, and $\infty$-norm of the vector $\bm{v}\in\mathbb{R}^{n}$, respectively. We use $\rho(\bm{M})$ to denote the spectral radius (i.e., the largest eigenvalue in absolute value) of a square matrix $\bm{M}\in\mathbb{R}^{n\times n}$. Given a set $\mathcal{X}$, $\Delta(\mathcal{X})$ denotes the set of probability distributions on $\mathcal{X}$. Let $x\sim P$ be a random variable with distribution $P\in\Delta(\mathcal{X})$. We use $\widehat{x}\sim P$ to denote a sample of the random variable $x$ drawn from the sample space $\mathcal{X}$ of $x$ according to the distribution $P$. We use $\bm{1}$, $\bm{0}$, and $\bm{I}$ to denote the all-one vector, the all-zero vector, and the identity matrix, respectively, with their dimension being clear from the context.
@@ 2. Optimal control of MDPs
$\mathbf{T}(\bm{v}^{\star})](s)\coloneqq\min_{a\in\mathcal{A}}\left\{\bm{c}(s,a)+\gamma\ \mathds{E}_{s^{+}\sim P(\cdot|s,a)}\left[\bm{v}^{\star}(s^{+})\right]\right\},\quad\forall s\in\mathcal{S}.$ (3)
@@
[…]
@@
[…]
@@ 4. Rank-one Q-learning (R1-QL)
where $\lambda_{k}\geq 0$ are properly chosen step-sizes satisfying the Robbins–Monro conditions (e.g., $\lambda_{k}=1/(k+1)$), and $\widehat{\mathbf{T}}_{k}$ is the empirical Bellman operator evaluated at iteration $k$, i.e., $\widehat{\mathbf{T}}_{k}(\bm{q}_{k})](s,a)\coloneqq[\widehat{\mathbf{T}}(\bm{q}_{k},\widehat{s}_{k}^{~+})](s,a)$ with $\widehat{s}_{k}^{~+}\sim P(\cdot|s,a)$ for each $(s,a)\in\mathcal{S}\times\mathcal{A}$ – the subscript $k$ in $\widehat{\mathbf{T}}_{k}$ denotes the dependence on the next-state sample $\widehat{s}_{k}^{~+}$ generated at iteration $k$.
@@
Now, using the approximation $\widehat{\bm{d}}_{k-1}\approx\widehat{\bm{P}}_{k-1}^{\top}\widehat{\bm{d}}_{k-1}$ (i.e., assuming $\widehat{\bm{d}}_{k-1}$ is a stationary distribution of $\widehat{\bm{P}}_{k-1}$ which does not hold exactly since $\widehat{\bm{d}}_{k-1}$ is only an approximation of a stationary distribution of $\widehat{\bm{P}}_{k-1}$), we derive
$\widehat{\bm{d}}_{k}=(1-\lambda_{k})\widehat{\bm{d}}_{k-1}+\lambda_{k}\bm{F}_{k}^{\top}\widehat{\bm{d}}_{k-1}.$
We note that the vector $\bm{f}=\bm{F}_{k}^{\top}\widehat{\bm{d}}_{k-1}$ can be computed using the following pseudo-code:
[…]
Observe that the approximation $\widehat{\bm{d}}_{k-1}\approx\widehat{\bm{P}}_{k-1}^{\top}\widehat{\bm{d}}_{k-1}$ significantly reduces the memory and time complexity of the algorithm since we do not need to keep track of the estimates $\widehat{\bm{P}}_{k}$ of the state-action transition probability matrix and perform full matrix-vector multiplications for updating the estimates $\widehat{\bm{d}}_{k}$ of the stationary distribution.
@@
8:    $\bm{f}(\widehat{s}_{k}^{~+},\widehat{a}_{k}^{~+})=\bm{f}(\widehat{s}_{k}^{~+},\widehat{a}_{k}^{~+})+\widehat{\bm{d}}_{k-1}(s,a);$
@@
10:   $\widehat{\bm{d}}_{k}=(1-\lambda_{k})\widehat{\bm{d}}_{k-1}+\lambda_{k}\bm{f}$; $\widehat{\bm{d}}_{k}=\widehat{\bm{d}}_{k}/\|\widehat{\bm{d}}_{k}\|_{1}$;
@@
The complete description of the proposed R1-QL algorithm is provided in Algorithm 2. We again note that the normalization is only introduced to avoid the accumulation of numerical errors. The following result discusses the convergence of the proposed algorithm.
@@ 6. Limitations and future research directions
where $e_{s}$ is the s{s}-th unit vector and $\mathbf{T}$ is now the undiscounted Bellman operator. Now, observe that
@@ Appendix A Technical Proofs
$\left\{\begin{aligned} &\bm{v}_{k+1}=\mathbf{T}(\bm{v}_{k})+\alpha_{k}\bm{1}\\ &\alpha_{k}=\frac{\gamma}{1-\gamma}\big\langle\bm{d}_{k},\mathbf{T}(\bm{v}_{k})-\bm{v}_{k}\big\rangle,\end{aligned}\right$. (19)
@@
$\left\{\begin{aligned} &\bm{v}_{k+1}=\mathbf{T}^{(k+1)}(\bm{v}_{0})+\beta_{k+1}\bm{1}\\ &\beta_{k+1}=\gamma\beta_{k}+\alpha_{k},\quad\text{\emph{with}}~\beta_{0}=0.\end{aligned}\right$.
@@
$\left\{\begin{aligned} &\bm{q}_{k+1}=(1-\lambda_{k})\bm{q}_{k}+\lambda_{k}\widehat{\mathbf{T}}_{k}(\bm{q}_{k})+\alpha_{k}\bm{1}\\ &\alpha_{k}=\lambda_{k}\Big(\frac{\gamma}{1-\gamma}\Big)\big\langle\widehat{\bm{d}}_{k},\widehat{\mathbf{T}}_{k}(\bm{q}_{k})-\bm{q}_{k}\big\rangle,\end{aligned}\right$. (23)
@@
$\left\{\begin{aligned} &\bm{q}_{k+1}=\bm{q}^{\emph{\text{QL}}}_{k+1}+\beta_{k+1}\bm{1}\\ &\beta_{k+1}=(1-\lambda_{k})\beta_{k}+\gamma\lambda_{k}\beta_{k}+\alpha_{k},\quad\text{\emph{with}}~\beta_{0}=0.\end{aligned}\right$. (24)
@@
[…]
@@
is a scaled, weighted average of Cesaro means of the sequences $\overline{\mathbf{T}}_{k}(\bm{q}^{\text{QL}}_{k})-\bm{q}^{\text{QL}}_{k}](s,a)$ that converge to zero almost surely for each $(s,a)\in\mathcal{S}\times\mathcal{A}$ (recall that $\lambda_{k}=1/(k+1)$ and $\bm{q}^{\text{QL}}_{k}\rightarrow\bm{q}^{\star}=\overline{\mathbf{T}}(\bm{q}^{\star})$ almost surely). ∎
@@ B.1. Algorithms
    $\begin{aligned} &\bm{z}_{k}=\bm{v}_{k}+\frac{1-\sqrt{1-\gamma^{2}}}{\gamma}(\bm{v}_{k}-\bm{v}_{k-1}),\\ &\bm{v}_{k+1}=\bm{z}_{k}+\frac{1}{1+\gamma}(\mathbf{T}(\bm{z}_{k})-\bm{z}_{k}).\end{aligned}$
@@
    $\begin{aligned} &\bm{z}_{k}=\bm{v}_{k}-\bm{v}_{k-1},\\ &\bm{z^{\prime}}_{k}=\mathbf{T}(\bm{v}_{k})-\mathbf{T}(\bm{v}_{k-1}),\\ &\delta_{k}=\begin{cases}0,&\bm{z}_{k}^{\top}(\bm{z}_{k}-\bm{z}_{k}^{\prime})=0,\\ \frac{\bm{z}_{k}^{\top}\big(\bm{v}_{k}-\mathbf{T}(\bm{v}_{k})\big)}{\bm{z}_{k}^{\top}(\bm{z}_{k}-\bm{z}_{k}^{\prime})},&\text{otherwise},\end{cases}\\ &\bm{v}_{k+1}=(1-\delta_{k})\mathbf{T}(\bm{v}_{k})+\delta_{k}\mathbf{T}(\bm{v}_{k-1}).\end{aligned}$
@@
    $\begin{aligned} &\texttt{for }(s,a)\in\mathcal{S}\times\mathcal{A}\\ &\qquad\widehat{s}^{~+}\sim P(\cdot|s,a),\\ &\qquad\bm{z}_{k}(s,a)=\bm{c}(s,a)+\gamma\min_{a^{+}\in\mathcal{A}}\bm{q}_{k}(\widehat{s}^{~+},a^{+}),\\ &\qquad\bm{z}_{k}^{\prime}(s,a)=\bm{c}(s,a)+\gamma\min_{a^{+}\in\mathcal{A}}\bm{q}_{k-1}(\widehat{s}^{~+},a^{+}),\\ &\texttt{endfor }\\ &\qquad\bm{q}_{k+1}=\bm{q}_{k}+\frac{1}{1+k}(\bm{z}_{k}^{\prime}-\bm{q}_{k})+\frac{k}{1+k}(\bm{z}_{k}-\bm{z}_{k}^{\prime}).\end{aligned}$
@@
    $\begin{aligned} &\bm{F}_{k}=\bm{0}\in\mathbb{R}^{|\mathcal{S}\times\mathcal{A}|\times|\mathcal{S}\times\mathcal{A}|},\\ &\texttt{for }(s,a)\in\mathcal{S}\times\mathcal{A}\\ &\qquad\widehat{s}^{~+}\sim P(\cdot|s,a),\\ &\qquad\widehat{a}^{~+}=\operatorname*{\arg\!\min}_{a^{+}\in\mathcal{A}}\bm{q}_{k}(\widehat{s}^{~+},a^{+}),\\ &\qquad\big[\widehat{\mathbf{T}}_{k}(\bm{q}_{k})\big](s,a)=\bm{c}(s,a)+\gamma\min_{a^{+}\in\mathcal{A}}\bm{q}_{k}(\widehat{s}^{~+},a^{+})-\bm{q}_{k}(s,a),\\ &\qquad\bm{F}_{k}\big((s,a),(\widehat{s}^{~+},\widehat{a}^{~+})\big)=1,\\ &\texttt{endfor }\\ &\widehat{\bm{P}}_{k}=\widehat{\bm{P}}_{k-1}+\frac{1}{2+k}(\bm{F}_{k}-\widehat{\bm{P}}_{k-1}),\\ &\bm{q}_{k+1}=\bm{q}_{k}+\frac{1}{1+k}(\bm{I}-\gamma\widehat{\bm{P}}_{k})^{-1}\big(\widehat{\mathbf{T}}_{k}(\bm{q}_{k})-\bm{q}_{k}\big).\end{aligned}$